# ComSD: Balancing Behavioral Quality and Diversity in Unsupervised Skill Discovery

### v1

@@ 3.2 Particle-based State Entropy Estimation
In decomposition Eq. 2, increasing the state entropy $H(\tau)$ encourages the agent to explore more widely and visit more state transitions. Following  Liu & Abbeel (2021a); Yarats et al. (2021b); Laskin et al. (2022b), we employ a particle-based entropy estimation proposed by  Liu & Abbeel (2021b). Concretely, $H(\tau)$ can be estimated by the Euclidean distance between each particle ($\tau$ in our paper) and its all $k$-nearest neighbor (Singh et al., 2003) in the latent space:
@@
$r^{intr}_{exploration}\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{batch}^{k-nn}(h_{i})}\log||h_{i}-h_{i}^{nn}||,\ \ where\ \ h_{i}=f_{\theta_{1}}(\tau_{i}).$ (10)
3.3 Skill-based Multi-objectives weighting
With the negative conditioned entropy $-H(\tau|z)$ estimated by $r^{intr}_{exploitation}$ and the state entropy estimated by $r^{intr}_{exploration}$, a naive way for intrinsic reward design is to employ a fixed coefficient $\alpha$ to scale the two terms ($r^{intr}_{naive}=r^{intr}_{exploration}+\alpha r^{intr}_{exploitation}$). However, we find it impossible to choose a proper fixed coefficient to balance them well especially in challenging multi-joint robot behavior discovery, which is because of a severe quality-versus-diversity conflict between the above two sub-intrinsic rewards. Specifically, when employing $r^{intr}_{exploration}$ alone, the agents can only learn dynamic behaviors of high activity but not easing movements or static postures, and the learned skills are highly homogeneous and indistinguishable by skill vectors (i.e., unsatisfactory exploitation). The naive intervention of $r^{intr}_{exploitation}$ can significantly improve the behavioral diversity, but it will simultaneously cause the lazy exploration, making the robots unable to discover active behaviors.
We propose Skill-based Multi-objectives Weighting (SMW), a simple but effective dynamic weighting mechanism for the issue above. SMW aims to take advantage of the two sub-rewards by setting different optimization objectives for different skill vectors. Concretely, it designs another changeable coefficient $\beta$ which is dynamically adjusted according to skill vectors. $\beta(z)$ is defined as:
@@ 3.3 Skill-based Multi-objectives weighting
The $\rm flag\it(\cdot)$ can be any function to map the high-dimensional skill vectors into 1-dimensional flag variables. We find that simply employing the first dimension of skill vectors (i.e., $\rm flag\it(z)=I_{1}^{T}\cdot z,\ {\rm where}\ I_{1}=(1,0,...,0)$) can perform well. The fixed hyper-parameters $w_{high},w_{low},f_{high},f_{low}$ and the clamp function $\rm clamp\it(\cdot)$ make the dynamic coefficient $\beta$ linearly related to flag variables $\rm flag\it(z)$ with different slopes in different regions. We clearly visualize the relation between $\beta$ and $\rm flag\it(z)$ in Figure 2. With SMW, the multi-objective intrinsic reward of ComSD is defined as:
$r^{intr}_{ComSD}=r^{intr}_{exploration}+\alpha\cdot\beta(z)\cdot r^{intr}_{exploitation}.$ (12)
Over the intrinsic-reward MDP $\mathcal{M}^{intr}$, the skill-conditioned policy $\pi(a|s,z)$ can be trained by RL with $z\sim p(z)$. Algorithm 1 in Appendix A provides the full pseudo-code of ComSD.
4 Skill Evaluation and Analysis
@@ 4.1 Experimental Setup
Environments. On the importance of environments: OpenAI Gym (Brockman et al., 2016) and DMControl (Tassa et al., 2018) are two most popular continuous robot locomotion benchmarks. In Gym, the episode is ended when agents lose their balance, while the episode length in DMControl is fixed. Laskin et al. (2022b) found this difference makes DMControl much harder for reward-free exploration since agents have to learn balance by themselves without any external signals.
@@
Following recent advanced methods (Laskin et al., 2022b; Yarats et al., 2021b; Liu & Abbeel, 2021a; b; Liu et al., 2023), we employ 16 downstream tasks of 4 domains from URLB (Laskin et al., 2021) and DMControl (Tassa et al., 2018) for skill evaluation. The domains Walker, Quadruped, Cheetah and Hopper are recognized as the most representative and challenging multi-joint robot environments, each of which contains 4 totally diverse and challenging downstream locomotion tasks.
Evaluations. In each domain, all the methods, including ComSD pre-train their agents for 2M environment steps with their respective intrinsic rewards. After unsupervised pre-training, the behavioral diversity and quality will be fully and reasonably evaluated on two adaptation tasks: skill combination (Eysenbach et al., 2018) and skill combination in URLB (Laskin et al., 2021) respectively, across all 16 downstream tasks with extrinsic reward. It means a total of 32 numerical results are employed for one method’s evaluation. DDPG (Lillicrap et al., 2015) is chosen as the backbone RL algorithm for all methods throughout the paper. The detailed settings of the skill combination and skill finetuning are provided along with the experimental results in Section 4.2.
Baselines. We compare our ComSD with five recent advanced skill discovery algorithms that are popular for robot behavior discovery. They are BeCL (Yang et al., 2023), CIC (Laskin et al., 2022b), APS (Liu & Abbeel, 2021a), SMM (Lee et al., 2019), and DIAYN (Eysenbach et al., 2018). All the methods try to optimize their MI target by designing intrinsic rewards for unsupervised RL. In addition to BeCL, they all try to maximize $I(\tau,z)$, where DIAYN and SMM choose the decomposition Eq. 1 while APS, CIC, and our ComSD choose decomposition Eq. 2. BeCL proposes a novel MI objective $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill. We provide a detailed description of these baselines in Appendix B.
The close baselines to our ComSD are CIC and APS. Apart from SMW, a novel weighting mechanism, ComSD still differs significantly from these methods. For CIC, ComSD follows it in state entropy estimation but first proposes to employ contrastive results for explicit state entropy maximization. APS and ComSD both explicitly optimize the state entropy and conditioned entropy, while ComSD utilizes contrastive learning for a more reasonable estimation. In addition, BeCL also employs contrastive learning like ComSD, but their contrastive components are different.
4.2 Main Results
Figure 3: The training curve of all 6 methods on 16 downstream skill combination tasks. ComSD outperforms all baselines significantly across all 16 tasks, demonstrating that ComSD discovers much more diverse and qualified behaviors than other methods for challenging multi-joint robots.
Skill Combination. In skill combination, we train another meta-controller $\pi^{\prime}(z|s)$ which selects the discovered skills automatically to achieve downstream tasks. The discovered skill-conditioned agent $\pi(a|s,z)$ is frozen, and $\pi^{\prime}(z|s)$ is trained by RL with extrinsic reward (see Section 2.1 and Appendix A.2 for a detailed problem definition). This adaptation task can effectively evaluate both the diversity and quality of the discovered skill set. For each method, 2M environment steps RL are allowed to train their meta-controllers. All the hyper-parameters of meta-controller and RL settings are shared across all the methods for a fair comparison, which we detail in Appendix C. The results of ComSD are obtained over 6 different random seeds.
The training curves are shown in Figure 3, which demonstrates that ComSD significantly outperforms all five baselines across all 16 downstream tasks. In skill combination, the better adaptation performance means that the behaviors discovered by ComSD are of higher diversity and quality than other methods. CIC and APS are the two most competitive baselines. Both of them and our ComSD employ the decomposition Eq. 2, which indicates the importance of explicitly maximizing state entropy for discovering robot behaviors. An interesting phenomenon is that CIC always gets a higher initial score than APS and ComSD but has no upward trend, which coincides with the fact that CIC can only find homogeneous behaviors at a high activity level. By contrast, APS and ComSD explicitly maximize the conditioned entropy through intrinsic rewards, greatly improving the diversity of learned skill sets and achieving higher final scores than CIC. Compared with APS, ComSD performs better due to the following two reasons: First, ComSD employs contrastive learning for a better estimation of both state entropy and conditioned entropy. Second, the proposed SMW enables ComSD to learn qualified behaviors at different activity levels, thus gaining better adaptation ability.
Skill Finetuning. In skill finetuning, a skill vector $z_{i}$ is chosen, and the corresponding skill $\pi(\cdot|s,z_{i})$ is finetuned with the extrinsic reward (see Section 2.1 and Appendix A.3 for a detailed problem definition). 100k environment steps with extrinsic reward are allowed for each method, where a proper skill vector is selected in the first 4k steps and the chosen skill is finetuned in another 96k steps. For skill selection, previous works employ random sampling (Yang et al., 2023), fixed choice (Laskin et al., 2022b; a), or reward-based choice (Liu & Abbeel, 2021a). We follow the official implementation of  Laskin et al. (2022b), employing a fixed mean skill with the first dimension set to 00. All the hyper-parameters of neural architecture and RL are shared across different methods for a fair comparison. We show the detailed experimental settings in Appendix C. The results of ComSD are obtained over 10 different random seeds.
Table 1: The numerical results of all 6 methods on 16 skill finetuning downstream tasks. ComSD exhibits competitive adaptation performance compared with state-of-the-art methods (CIC and BeCL). In skill finetuning, only one skill is used for evaluation. The behavioral diversity is also ignored. It is one-sided to employ skill finetuning alone for evaluation, like recent works.
@@ 4.2 Main Results
The numerical results on 16 downstream tasks are shown in Table 1. CIC and BeCL are the state-of-the-art methods for skill finetuning. ComSD performs comparable or better on 10/16 tasks than CIC and outperforms BeCL across 15/16 tasks, also achieving state-of-the-art adaptation performance. Actually, the score of the initial few steps has little to do with the final score in state-based locomotion RL, which means it’s hard to find a proper skill from a large skill set within only 4k steps. This explains why CIC and BeCL use random or fixed choices. Furthermore, higher behavioral diversity causes higher choice difficulty in ComSD. ComSD discovers the most diverse behaviors (state-of-the-art adaptation on skill combination) while simultaneously achieving competitive performance on skill finetuning with a fixed skill, which further shows its superior behavioral quality.
@@
Figure 4: Numerical ablation experiments on (left) two skill combination tasks and (right) two skill finetuning tasks. Contrastive estimation of negative conditioned entropy and SMW are both necessary for ComSD to achieve the most competitive results on both kinds of adaptation tasks.
Figure 5: Skill activity analysis of our ComSD, ComSD w/o SMW, and all baselines. We estimate the skill activity by visited state entropy (KTD and AKD) and report the activity variance and range of uniformly selected walker skills for each method. ComSD exhibits much higher entropy variance and range, demonstrating that ComSD indeed discovers skills at different activity levels.
4.3 Numerical Ablation Study
In this section, we conduct numerical ablation experiments to show the effectiveness of both contrastive conditioned entropy estimation and SMW. The results are shown in Figure 4, where CIC is actually ComSD w/o explicit conditioned entropy estimation&SMW, and APS is actually ComSD w/o contrastive conditioned entropy estimation&SMW. Compared with implicit maximization (CIC), explicitly maximizing the skill-conditioned entropy (APS and ComSD w/o SMW) improves the training curve slope on skill combination, which means better exploitation, i.e., behavioral diversity. Moreover, the contrastive result is a better conditioned entropy estimation, improving APS on both kinds of adaptation tasks. However, simply maximizing the conditioned entropy explicitly leads to a trivial solution in practice, where the agent gives up continuous movement and exploration. They can only learn static postures but not dynamic behaviors, exhibiting performance drops on skill finetuning compared with CIC. This phenomenon is also observed by Laskin et al. (2022b). SMW effectively alleviates this exploration hurt by setting multiple optimization objectives. It improves behavioral exploitation while maintaining the tendency toward exploration, making ComSD exhibit state-of-the-art performance on both kinds of adaptation tasks. In summary, SMW and contrastive conditioned entropy estimation are both necessary for ComSD.
4.4 Quality Analysis of Learned Skills
Numerical results on 32 downstream tasks have demonstrated the effectiveness and superiority of ComSD. In this section, we directly analyze the learned skills to show that our method indeed discovers behaviors at different activity levels, which previous advanced methods cannot. For each method, we uniformly select eleven walker skills and use the entropy of visited states to represent the skill activity. K-Th-nearest-neighbor Distance (KTD) and All-K-nearest-neighbor Distance (AKD) (Liu & Abbeel, 2021b), are employed as state entropy estimations. KTD computes the k-th-nearest-neighbor Euclidean distance of each state and takes the mean, while AKD considers all the k-nearest neighbors of each state and takes the mean distance. For each method, we report the entropy variance and entropy range of the selected skills in Figure 5. ComSD discovers a skill set with the biggest activity range and variance, demonstrating it indeed produces skills at diverse activity levels, where mainly attributes to multiple optimization targets brought by multi-objective intrinsic reward. We strongly refer readers to Appendix D for additional visualization and analysis.

### v2

@@ 3.2 Particle-based State Entropy Estimation
In decomposition Eq. (2), increasing the state entropy $H(\tau)$ encourages the agent to explore more widely and visit more state transitions. Following  Liu and Abbeel (2021a); Yarats et al. (2021b); Laskin et al. (2022b), we employ a particle-based entropy estimation proposed by  Liu and Abbeel (2021b). Concretely, $H(\tau)$ can be estimated by the Euclidean distance between each particle ($\tau$ in our paper) and its all $k$-nearest neighbor Singh et al. (2003) in the latent space:
@@
$r_{exploration}\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{batch}^{k-nn}(h_{i})}\log||h_{i}$ $-h_{i}^{nn}||,$ (10)
$where\ \ h_{i}$ $=f_{\theta_{1}}(\tau_{i}).$
Figure 3: The training curve of all 6 methods on 16 downstream skill combination tasks. ComSD outperforms all baselines significantly across all 16 tasks, demonstrating that ComSD discovers much more exploratory and diverse behaviors than other methods for challenging multi-joint robots.
3.3 Skill-based Multi-objective Weighting
With the negative conditioned entropy $-H(\tau|z)$ estimated by $r^{contrast}_{diversity}$ and the state entropy estimated by $r_{exploration}$, a naive way for intrinsic reward design is to employ a fixed coefficient $\alpha$ to scale the two terms ($r^{intr}_{naive}=r_{exploration}+\alpha\cdot r^{contrast}_{diversity}$). However, in practice, we find it impossible to use a fixed coefficient to balance them well when potential behaviors are complex and hard to distinguish by only ending states. Concretely, when employing $r_{exploration}$ alone, the robots can only learn dynamic behaviors of high activity but not other exploratory levels, and the learned behaviors are highly homogeneous. Increasing the fixed weight $\alpha$ of $r^{contrast}_{diversity}$ (or other $-H(\tau|z)$ estimation in other methods) can surely improve the behavioral diversity, but it will simultaneously reduce the exploration (see experiments in Figure 6), making robots prefer static postures (distinguishable by ending states) and ignore active behaviors (indistinguishable by ending states). This is because distilling ending states into skill vectors is an easy-to-reach local optimal solution for state-skill MI maximization.
We propose Skill-based Multi-objective Weighting (SMW), an easy-but-effective weighting mechanism for the issue above. The main idea of SMW is to set different learning objectives for different skills instead of using fixed learning objectives across all skills. Concretely, it designs another changeable coefficient $\beta(z)$ which is dynamically adjusted according to skill vectors. $\beta(z)$ is defined as:
@@ 3.3 Skill-based Multi-objective Weighting
[…]
$r^{intr}_{ComSD}=r_{exploration}+\alpha\cdot\beta(z)\cdot r^{contrast}_{diversity}.$ (12)
Over the intrinsic-reward MDP $\mathcal{M}^{intr}$ augmented by $r^{intr}_{ComSD}$, ComSD trains the skill-conditioned policy $\pi(a|s,z)$ via RL with $z$ sampled by uniform distribution. Appendix A provides the full pseudo-code of ComSD.
4 Experiments: Skill Adaptation and Analysis
@@ 4.1 Experimental Setup
Environments. On the importance of environments: OpenAI Gym Brockman et al. (2016) and DMC Tassa et al. (2018) are two most popular continuous robot locomotion benchmarks. In Gym, the episode is ended when agents lose their balance, while the episode length in DMC is fixed. Laskin et al. (2022b) found this difference makes DMC much harder for reward-free exploration since agents have to learn balance by themselves without any external signals.
@@
Following recent advanced methods Laskin et al. (2022b); Yarats et al. (2021b); Liu et al. (2023); Laskin et al. (2020); Stooke et al. (2021); Yarats et al. (2021a), we employ 16 downstream tasks of 4 domains from URLB Laskin et al. (2021) and DMC Tassa et al. (2018) for skill adaptation evaluation. The domains Walker, Quadruped, Cheetah, and Hopper are recognized as the most representative and challenging multi-joint robot environments, each of which contains four totally diverse and challenging downstream tasks.
@@
Table 1: The numerical results of all 6 methods on 16 skill finetuning downstream tasks. ComSD exhibits competitive adaptation performance compared with state-of-the-art methods (CIC and BeCL).
Evaluations. In each domain, all the methods pre-train their agents for 2M environment steps with their respective intrinsic rewards. After unsupervised pre-training, the behavioral quality will be fully evaluated on two adaptation tasks: skill combination Eysenbach et al. (2018) and skill combination in URLB Laskin et al. (2021) respectively, across all 16 downstream tasks with extrinsic reward. It means a total of 32 numerical results are employed for one method’s evaluation. DDPG Lillicrap et al. (2015) is chosen as the backbone RL algorithm for all methods throughout the paper. The detailed settings of the skill combination and skill finetuning are provided along with the experimental results in Section 4.2. The direct quality analysis is also provided in Section 4.3.
Baselines. We compare our ComSD with five recent advanced algorithms for robot behavior discovery. They are BeCL Yang et al. (2023), CIC Laskin et al. (2022b), APS Liu and Abbeel (2021a), SMM Lee et al. (2019), and DIAYN Eysenbach et al. (2018). We provide detailed description of these baselines in Appendix B. The close baselines to our ComSD are CIC and APS, where they all design intrinsic rewards based on Eq. (2). Apart from the most notable difference, a novel weighting mechanism (SMW), ComSD still significantly differs from these methods. CIC doesn’t design a diversity reward, while ComSD proposes a novel contrastive diversity reward to explicitly encourage diversity. APS also employs both diversity reward and exploration reward, but it doesn’t conduct contrastive learning which is the key to provide effective incentives for complex agents. In addition, BeCL also employs contrastive learning like us, but the contrastive components are totally different.
@@
Figure 4: Adaptation ablation experiments on (left) two skill combination tasks and (right) two skill finetuning tasks. Contrastive diversity reward and SMW are both necessary for ComSD to achieve the most competitive results on both kinds of adaptation tasks.
Figure 5: Left three: Behavioral exploration analysis of all methods. AKD can represent the exploratory level of one skill. ComSD leads on all three metrics, which means it generates behaviors at different exploratory levels, including dynamic movements and static postures, which other methods cannot. Right one: Behavioral diversity analysis of all methods. ComSD discovers the most diverse robot behaviors.
4.2 Adaptation Results
Skill Combination. In skill combination, we train another meta-controller $\pi^{\prime}(z|s)$ which selects the discovered skills automatically to achieve downstream tasks. The discovered skill-conditioned agent $\pi(a|s,z)$ is frozen, and $\pi^{\prime}(z|s)$ is trained by RL with extrinsic reward (see Section 2.1 and Appendix A for a detailed problem definition). This adaptation task can effectively evaluate behavioral quality (both exploration and diversity). For each method, 2M environment steps RL are allowed to train their meta-controllers. All the hyper-parameters are shared across all the methods for a fair comparison, which we detail in Appendix C. The results of ComSD are obtained over 6 different random seeds.
The training curves are shown in Figure 3, which demonstrates that ComSD significantly outperforms all five baselines across all 16 downstream tasks. In skill combination, the better adaptation performance means that the behaviors discovered by ComSD are of higher quality (both exploration and diversity) than other methods. CIC and APS are the two most competitive baselines. Both of them and our ComSD employ the decomposition Eq. (2), which indicates the importance of explicitly maximize exploration. An interesting phenomenon is that CIC always gets a higher initial score than APS and ComSD but has no upward trend, which coincides with the fact that CIC can only find homogeneous behaviors at a high exploratory level. By contrast, APS and ComSD explicitly maximize the conditioned entropy through diversity reward, greatly improving the diversity of learned skill sets and achieving higher final scores than CIC. Compared with APS, ComSD performs better due to the following two reasons: First, the contrastive learning employed by ComSD provides stronger incentives for skill exploration and exploitation in complex domains. Second, the proposed SMW enables ComSD to learn diverse behaviors at different activity levels, thus gaining better multi-task adaptation.
Skill Finetuning In skill finetuning, a skill vector $z_{i}$ is chosen, and the corresponding skill $\pi(\cdot|s,z_{i})$ is finetuned with the extrinsic reward (see Section 2.1 and Appendix A for a detailed problem definition). 100k environment steps with extrinsic reward are allowed for each method, where one skill is selected in the first 4k steps and then finetuned in another 96k steps. For skill selection, previous works employ random sampling Yang et al. (2023), fixed choice Laskin et al. (2022b, a), or reward-based choice Liu and Abbeel (2021a). We follow the official implementation of  Laskin et al. (2022b), employing a fixed mean skill with the first dimension set to 00. All the hyper-parameters are shared across different methods for a fair comparison. We show the detailed experimental settings in Appendix C. The results of ComSD are obtained over 10 different random seeds.
The numerical results on 16 downstream tasks are shown in Table 1. CIC and BeCL are the state-of-the-art methods for skill finetuning. ComSD performs comparable or better on 10/16 tasks than CIC and outperforms BeCL across 15/16 tasks, also achieving state-of-the-art adaptation performance. Actually, the score of the initial few steps has little to do with the final score in state-based locomotion RL, which means it’s hard to find a proper skill from a large skill set within only 4k steps. This explains why CIC and BeCL use random or fixed choices. Furthermore, higher behavioral diversity causes higher choice difficulties. ComSD discovers the most diverse behaviors (see skill combination results and analysis in section 4.3) while simultaneously achieving competitive performance on skill finetuning, which further shows its superior skill quality.
Adaptation Ablation. We conduct ablation experiments on both kinds of adaptation tasks to show the effectiveness of (i) contrastive diversity reward $r^{contrast}_{diversity}$ and (ii) SMW in Figure 4, where CIC is actually ComSD w/o $r^{contrast}_{diversity}$&SMW, and APS is actually ComSD w/o contrastive learning&SMW. Compared with no diversity reward (CIC), designing diversity reward (APS and ComSD w/o SMW) enables better behavioral diversity, corresponding to larger curve slope on skill combination. Moreover, contrastive learning enables ComSD w/o SMW to outperform APS on both kinds of adaptation tasks. In practice, a fixed diversity reward (APS and ComSD w/o SMW) reduces the behavioral exploration (see Section 4.3). This causes severe performance drops on skill finetuning, which is effectively alleviated by SMW. In summary, both $r^{contrast}_{diversity}$ and SMW are needed for ideal adaptation.
4.3 Behavioral Quality Analysis
Behavioral Exploration. In this section, we analyze the behavioral exploration (activity) of all methods. We use the visited-state coverage (entropy) of one skill to represent its exploration. The visited state entropy is estimated by the All-K-nearest-neighbor Distance (AKD) Liu and Abbeel (2021b) that computes the Euclidean distance between each state and its k nearest neighbors, and then takes the mean. AKD can denote the exploratory level of one skill. For each method, we use grid sweep  Laskin et al. (2022b) to select eleven walker skills and report the variance, range, and maximum of AKD in Figure 5. ComSD and CIC lead on AKD-max, which means they both discover the most exploratory skill while other methods cannot. In addition, ComSD exhibits huge superiority in skill activity range and variance, indeed discovering behaviors at different exploratory levels, which all other methods cannot.
Behavioral Diversity. In this section, we try to analyze the behavioral diversity of all methods. For one skill, we represent it by the Mean State (MS) over 1k visited states. For each method, we use grid sweep  Laskin et al. (2022b) to select several walker behaviors, and then calculate the MS-coverage by k-nearest-neighbor Liu and Abbeel (2021b) Euclidean distance. MS-coverage can partially represent behavioral diversity to some extent. The results in Figure 5 demonstrate ComSD has huge advantages on MS-coverage, discovering the most diverse skill set.
SMW Moltivation & Behavioral Quality Ablation. We ablate SMW, setting different fixed coefficients $\alpha$ of $r^{contrast}_{diversity}$ to show the effect on behavioral quality. As shown in Figure 6 left, there exists a serious conflict between exploration and diversity. Concretely, $r^{contrast}_{diversity}$ simultaneously improves behavioral diversity and hurts exploration, and so do other diversity incentives in other methods (e.g., APS). Any fixed weight $\alpha$ cannot achieve diversity-exploration balance. This motivates our SMW which links the weight to the changing skill vectors, achieving dynamic weighting, as mentioned in Section 3.3. In addition, SMW further enables behavior discovery at different exploratory levels (Figure 6 right), which a fixed coefficient also cannot support.
Skill Visualization. We visualize the behaviors of ComSD, CIC, and APS in Appendix D, proving that (i) CIC can only learn dynamic but homogeneous flipping behaviors; (ii) APS acquires different static postures of lazy exploration; and (iii) ComSD discovers diverse behaviors at different exploratory levels, including flipping, standing, falling down, and different yoga postures. ComSD’s qualified behaviors enable state-of-the-art performance on different adaptation tasks.
Figure 6: SMW motivation & Behavioral quality ablation. Left: The exploration-diversity conflict in complex domains. Increasing the fixed weight $\alpha$ of diversity reward can surely improve the behavioral diversity, but it will simultaneously reduce the skill exploration. The dynamic SMW can effectively achieve diversity-exploration balance. Right: SMW enables behaviors at different exploratory levels, further improving behavioral quality.